40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Jan 4 19:04:29 2019, awade, DailyProgress, RFAM, Time to look at RF AM and RIN 
    Reply  Wed Jan 9 17:25:38 2019, anchal, DailyProgress, RFAM, Time to look at RF AM and RIN RIN_DBN_Coherence_Measurement.pdf
       Reply  Wed Jan 9 19:05:56 2019, anchal, DailyProgress, RFAM, Time to look at RF AM and RIN 
          Reply  Thu Jan 10 11:38:53 2019, awade, DailyProgress, RFAM, Time to look at RF AM and RIN 
             Reply  Thu Jan 10 17:30:24 2019, anchal, DailyProgress, RFAM, Time to look at RF AM and RIN 
                Reply  Thu Jan 10 18:52:20 2019, anchal, DailyProgress, RFAM, Time to look at RF AM and RIN 
                   Reply  Sat Jan 12 13:11:02 2019, awade, DailyProgress, RFAM, Time to look at beatnote after RF AM and RIN tuneup 
                      Reply  Wed Jan 16 11:07:41 2019, anchal, DailyProgress, NoiseBudget, Beatnote after RF AM and RIN tuneup LaserPowerNoise_to_BNFreqNoise_TF.pdfIN_to_BN_TF.zipTransmission_RIN_Jan15_2019.pdfRIN_Measurement_20190115.zip20190116_130937noiseBudget.pdf
    Reply  Thu Jan 10 14:19:46 2019, anchal, DailyProgress, RFAM, Time to look at RF AM and RIN 
Message ID: 2276     Entry time: Wed Jan 9 19:05:56 2019     In reply to: 2275     Reply to this: 2278
Author: anchal 
Type: DailyProgress 
Category: RFAM 
Subject: Time to look at RF AM and RIN 

Preliminary testing

I hooked up an HP8560E to RF port of the South FSS Refl RFPD. I unlocked laser from the cavity and brought it to a place which seemed far away from any mode of the cavity.

I was hoping to see a peak at 37 MHz which is PDH modulation frequency of FSS on this path. To my surprise, signal level was ~-80 dBm only at 37 MHz. On zooming out, I saw a 29.5 MHz peak instead of about ~-55dBm. @9.5 MHz is second harmonic of the modulation frequency for South PMC PDH loop. This seemed weird as sidebands of EOM before PMC should not reach after PMC. So I have taken noted some preliminary numbers for future. This isn't complete analysis:

Laser Slow Freq Offset (V) Peak Frequency (MHz) Peak Value (dBm)
-5.9975 0 -11
-5.9975 29.5 -55
-5.9975 37 -60
-5.9875 0 -11
-5.9875 29.5 -55
-5.9875 37 -60
-6.0275 0 -11
-6.0275 29.5 -55
-6.0275 37 -70

All data was taken with a span of 50 MHz. Note how on changing the frequency offset, I got a offset point where 37 MHz Peak almost disappears but the 29.5 MHz peak is present irrespective of the frequency. Need to investigate more.


Thu Jan 10 14:31:56 2019, in reply 2278

It is a resonant EOM.

ELOG V3.1.3-